OpenStudy (anonymous):

What equation results from completing the square and then factoring? x^2 + 4x = 12 A. (x + 2)2 = 8 B. (x + 2)2 = 10 C. (x + 2)2 = 16 D. (x + 2)2 = 14

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

no i dont know how do i solve for this?

OpenStudy (anonymous):

Well for completing the you take one half of the coefficient of x, square it and then add it to both sides, so the "4x" (4/2) = 2^2 = 4 \[x^2+4x+4=12+4 \implies x^2+4x+4=16\] Now you can facto the left side to get \[(x+2)^2=16\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

ok thank you big help my quiz is all questions like that :) really helps

OpenStudy (anonymous):

I did not know it was a quiz, we are not allowed to help on quiz's as that's for you to test what you've learned but since you already know the steps, I prefer you go over them yourself and do the same with the rest, and not cheat. Good luck.

OpenStudy (anonymous):

study guide for a quiz we get a exact copy of the quiz to study

OpenStudy (anonymous):

Oh ok, in that case, I hope that helped! Good luck on your quiz friend!

OpenStudy (anonymous):

it did and i really appreciate it

OpenStudy (anonymous):

No problem ^.^